40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 8 of 327  Not logged in ELOG logo
ID Date Author Type Category Subject
  16135   Wed May 12 14:23:20 2021 JordanUpdateSUSMass Properties of SOS Assembly with 3"->2" Optic sleeve, in SI units
Attachment 1: Moments_of_Inertia_SI.PNG
Moments_of_Inertia_SI.PNG
  16134   Wed May 12 13:06:15 2021 Ian MacMillanUpdateCDSSUS simPlant model

Working with Chris, we decided that it is probably better to use a simple filter module as a controller before we make the model more complicated. I will use the plant model that I have already made (see attachment 1 of this). then attach a single control filter module to that: as seen in attachment 1. because I only want to work with one degree of freedom (position) I will average the four outputs which should give me the position. Then by feeding the same signal to all four inputs I should isolate one degree of freedom while still using the premade plant model.

The model I made that is shown in attachment 2 is the model I made from the plan. And it complies! yay! I think there is a better way to do the average than the way I showed. And since the model is feeding back on itself I think I need to add a delay which Rana noted a while ago. I think it was a UnitDelay (see page 41 of RTS Developer’s Guide). So I will add that if we run into problems but I think there is enough going on that it might already be delayed.

Since our model (x1sup_isolated.mdl) has compiled we can open the medm screens for it. I provide a procedure below which is based on Jon's post

[First start the cymac and have the model running]
$  cd docker-cymac
$  eval $(./env_cymac)

$  medm -x /opt/rtcds/tst/x1/medm/x1sup_isolated/X1SUP_ISOLATED_GDS_TP.adl

To see a list of all medm screens use:

$  cd docker-cymac
$  ./login_cymac
 #  cd /opt/rtcds/tst/x1/medm/x1sup_isolated
 #  ls

Some of the other useful ones are:

adl screen Description
X1SUP_ISOLATED_Control_Module.adl This is the control filter module shown in attachment 2 at the top in the center. This module will represent the control system.
X1SUP_ISOLATED_C1_SUS_SINGLE_PLANT_Plant_POS_Mod.adl

See attachment 4. This screen shows the POS plant filter module that will be filled by the filter representing the transfer function of a damped harmonic oscillator:        \frac{x}{F}=\frac{\omega_0^2}{\omega_0^2+i\frac{\omega_0 \omega}{Q}-\omega^2}

THIS TF HAS BEEN UPDATED SEE NEXT POST

The first one of these screens that are of interest to us (shown in attachment 3) is the X1SUP_ISOLATED_GDS_TP.adl screen, which is the CDS runtime diagnostics screen. This screen tells us "the success/fail state of the model and all its dependencies." I am still figuring out these screens and the best guide is T1100625.

The next step is taking some data and seeing if I can see the position damp over time. To do this I need to:

  1. Edit the plant filter for the model and add the correct filter.
  2. Figure out a filter for the control system and add it to that. (I can leave it as is to see what the plant is doing) 
  3. Take some position data to show that the plant is a harmonic oscillator and is damping away.
Attachment 1: SimplePlant_SingleContr.pdf
SimplePlant_SingleContr.pdf
Attachment 2: x1sup_isolated.pdf
x1sup_isolated.pdf
Attachment 3: X1SUP_ISOLATED_GDS_TP.png
X1SUP_ISOLATED_GDS_TP.png
Attachment 4: X1SUP_ISOLATED_C1_SUS_SINGLE_PLANT_Plant_POS_Mod.png
X1SUP_ISOLATED_C1_SUS_SINGLE_PLANT_Plant_POS_Mod.png
  16133   Wed May 12 11:45:13 2021 Anchal, PacoSummarySUSNew IMC Settings are miserable

We picked a few parameters from 40m summary page and plotted them to see the effect of new settings. On April 4th, old settings were present. On April 28th (16091), new input matrices and F2A filters were uploaded but suspension gains remained the same. On May 5th (16120), we uploaded new (higher) suspension gains. We chose Sundays on UTC so that it lies on weekends for us. Most probably nobody entered 40m and it was calmer in the institute as well.

  • On MC_F spectrum, we see that that noise decreased in 0.3-0.7 Hz but there is more noise from 1-1.5 Hz.
  • On MC_TRANS_QPD, we see that both TRANS PIT and YAW signals were almost twice as noisy.
  • On MC_REFL_DC too, we see that the noise during the locked state seems to be higher in the new configuration.

We can download data and plot comparisons ourselves and maybe calculate the spectrums of MC_TRANS_PIT/YAW and MC_REFL_DC when IMC was locked. But we want to know if anyone has better ways of characterizing the settings that we should know of before we get into this large data handling which might be time-consuming. From this preliminary 40m summary page plots, maybe it is already clear that we should go back to old settings. Awaiting orders.

 

Attachment 1: MC_F_Comparison.pdf
MC_F_Comparison.pdf
Attachment 2: MC_TRANS_QPD_Comparison.pdf
MC_TRANS_QPD_Comparison.pdf
Attachment 3: IMC_REFL_DC_Comparison.pdf
IMC_REFL_DC_Comparison.pdf
  16132   Wed May 12 10:53:20 2021 Anchal, PacoUpdateLSCPSL-IMC PDH Loop and XARM PDH Loop diagram

Attached is the control loop diagram when main laser is locked to IMC and a single arm (XARM) is locked to the transmitted light from IMC.

Quote:
 
  • I'll post a clean loop diagram soon to make this loopology clearer.

 

Attachment 1: IMC_SingleArm.pdf
IMC_SingleArm.pdf
  16131   Tue May 11 17:43:09 2021 KojiUpdateCDSI/O Chassis Assembly

Did you match the local PC time with the GPS time?

  16130   Tue May 11 16:29:55 2021 JonUpdateCDSI/O Chassis Assembly
Quote:

Timing system set-up

The next step is to provide the 65 kHz clock signals from the timing fanout via LC optical fiber. I overlooked the fact that an SPX optical transceiver is required to interface the fiber to the timing slave board. These were not provided with the timing slaves we received. The timing slaves require a particular type of transceiver, 100base-FX/OC-3, which we did not have on hand. (For future reference, there is a handy list of compatible transceivers in E080541, p. 14.) I placed a Digikey order for two Finisar FTLF1217P2BTL, which should arrive within two days.

Today I brought and installed the new optical transceivers (Finisar FTLF1217P2BTL) for the two timing slaves. The timing slaves appear to phase-lock to the clocking signal from the master fanout. A few seconds after each timing slave is powered on, its status LED begins steadily blinking at 1 Hz, just as in the existing 40m systems.

However, some other timing issue remains unresolved. When the IOP model is started (on either FE), the DACKILL watchdog appears to start in a tripped state. Then after a few minutes of running, the TIM and ADC indicators go down as well. This makes me suspect the sample clocks are not really phase-locked. However, the models do start up with no error messages. Will continue to debug...

Attachment 1: Screen_Shot_2021-05-11_at_3.03.42_PM.png
Screen_Shot_2021-05-11_at_3.03.42_PM.png
  16129   Mon May 10 18:19:12 2021 Anchal, PacoUpdateLSCIMC WFS noise contribution in arm cavity length noise, Corrections

A few corrections to last analysis:

  • The first plot was not IMC frequency noise but actually MC_F noise budget.
    • MC_F is frequency noise in the IMC FSS loop just before the error point where IMC length and laser frequency is compared.
    • So, MC_F (in high loop gain frequency region upto 10kHz) is simply the quadrature noise sum of free running laser noise and IMC length noise.
    • Between 1Hz to 100 Hz, normally MC_F is dominated by free running laser noise but when we injected enough angular noise in WFS loops, due to Angle to length coupling, it made IMC length noise large enough in 25-30 Hz band that we started seeing a bump in MC_F.
    • So this bump in MC_F is mostly the noise due to Angle to length coupling and hence can be used to calculate how much Angular noise normally goes into length noise.
  • In the remaining plots, MC_F was plotted with conversion into arm length units but this was wrong. MC_F gets suppressed by IMC FSS open loop gain before reaching to arm cavities and hence is hardly present there.
  • The IMC length noise however is not suppresed until after the error point in the loop. So the length noise (in units of Hz calculated in the first step above) travels through the arm cavity loop.
  • We already measured the transfer function from ITMX length actuation to XARM OUT, so we know how this length noise shows up at XARM OUT.
  • So in the remaining plots, we plot contribution of IMC angular noise in the arm cavities. Note that the factor of 3 business still needed to be done to match the appearance of noise in XARM_OUT and YARM_OUT signal from the IMC angular noise injection.
  • I'll post a clean loop diagram soon to make this loopology clearer.
Attachment 1: ArmCavNoiseContributions.pdf
ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf
  16128   Mon May 10 10:57:54 2021 Anchal, PacoSummaryCalibrationUsing ALS beatnote for calibration, test

Test details:

  • We locked both arms and opened the shutter for Yend green laser.
  • After toggling the shutter on.off, we got a TEM00 mode of green laser locked to YARM.
  • We then cleared the phase Y history by clicking "CLEAR PHASE Y HISTROY" on C1LSC_ALS.adl (opened from sitemap > ALS > ALS).
  • We sent excitation signal at ITMY_LSC_EXC using awggui at 43Hz, 77Hz and 57Hz.
  • We measured the power spectrum and coherence of C1:ALS-BEATY_FINE_PHASE_OUT_HZ_DQ and C1:SUS-ITMY_LSC_OUT_DQ.
  • The BEATY_FINE_PHASE_OUT_HZ is already calibrated in Hz. This we assume is done by multip[lying the VCO slope in Hz/cts to the error signal of the digital PLL loop that tracks the phase of beatnote.
  • We calibrated C1:SUS-ITMY_LSC_OUT_DQ by multiplying with
    \large 3 \times \frac{2.44 \, nm/cts}{f^2} \times \frac{c}{1064\,nm \times 37.79\, m} = \frac{54.77}{f^2} kHz/cts where f is in Hz.
    The 2.44/f2 nm/cts is taken from 13984.
  • We added the calibration as Poles/zeros option in diaggui using gain=54.577e3 and poles as "0, 0".
  • We found that ITMY_LSC_OUT_DQ calibration matches well at 57Hz but overshoots (80 vs 40) at 43 Hz and undershoots (50 vs 80) at 77Hz.

Conclusions:

  • If we had DRFPMI locked, we could have used the beatnote spectrum as independent measurement of arm lengths to calibrate the interferometer output.
  • We can also use the beatnote to confirm or correct the ITM actuator calibrations. Maybe shape is not exactly 1/f2 unless we did something wrong here or the PLL bandwidth is too short.
Attachment 1: BeatY_ITMY_CalibrationAt57Hz.pdf
BeatY_ITMY_CalibrationAt57Hz.pdf BeatY_ITMY_CalibrationAt57Hz.pdf
Attachment 2: BeatY_ITMY_CalibrationAt43Hz.pdf
BeatY_ITMY_CalibrationAt43Hz.pdf BeatY_ITMY_CalibrationAt43Hz.pdf
Attachment 3: BeatY_ITMY_CalibrationAt77Hz.pdf
BeatY_ITMY_CalibrationAt77Hz.pdf BeatY_ITMY_CalibrationAt77Hz.pdf
  16127   Fri May 7 11:54:02 2021 Anchal, PacoUpdateLSCIMC WFS noise contribution in arm cavity length noise

We today measured the calibration factors for XARM_OUT and YARM_OUT in nm/cts and replotted our results from 16117 with the correct frequency dependence.


Calibration of XARM_OUT and YARM_OUT

  • We took transfer function measurement between ITMX/Y_LSC_OUT and X/YARM_OUT. See attachment 1 and 2
  • For ITMX/Y_LSC_OUT we took calibration factor of 3*2.44/f2 nm/cts from 13984. Note that we used the factor of 3 here as Gautum has explicitly written that the calibration cts are DAC cts at COIL outputs and there is a digital gain of 3 applied at all coil output gains in ITMX and ITMY that we confirmed.
  • This gave us callibration factors of XARM_OUT: 1.724/f2 nm/cts , and YARM_OUT: 4.901/f2 nm/cts. Note the frrequency dependence here.
  • We used the region from 70-80 Hz for calculating the calibration factor as it showed the most coherence in measurement.

Inferring noise contributions to arm cavities:

  • For converting IMC frequency noise to length noise, we used conversion factor given by \lambda L / c where L is 37.79m and lambda is wavelength of light.
  • For converting MC1 ASCPIT OUT cts data to frequency noise contributed to IMC, we sent 100,000 amplitude bandlimited noise  from 25 Hz to 30 Hz at C1:IOO-MC1_PIT_EXC. This noise was seen at both MC_F and ETMX/Y_LSC_OUT channels. We used the noise level at 29 Hz to get a calibration for MC1_ASCPIT_OUT to IMC Frequency in Hz/cts. This measurement was done in 16117.
  • Once we got the calibration above, we measured MC1_ASCPIT_OUT power spectrum without any excitaiton and multiplied it with the calibration factor.
  • Attachment 3 is our main result.
    • Page 1 shows the calculation of Angle to Length coupling by reading off noise injects in MC1_ASCPIT_OUT in MC_F. This came out to 10.906/f2 kHz/cts.
    • Page 2-3 show the injected noise in X arm cavity length units. Page 3 is the zoomed version to show the matching of the 2 different routes of calibration.
    • BUT, we needed to remove that factor of 3 we incorporated earlier to make them match.
    • Page 4 shows the noise contribution of IMC angular noise in XARM cavity.
    • Page 5-6 is similar to 2-3 but for YARM. The red note above applied here too! So the factor of 3 needed to be removed in both places.
    • Page 7 shows the noise contribution of IMC angular noise in XARM cavity.

Conclusions:

  • IMC Angular noise contribution to arm cavities is atleast 3 orders of magnitude lower then total armc cavity noise measured.

Edit Mon May 10 18:31:52 2021

See corrections in 16129.

Attachment 1: ITMX-XARM_TF.pdf
ITMX-XARM_TF.pdf ITMX-XARM_TF.pdf
Attachment 2: ITMY-YARM_TF.pdf
ITMY-YARM_TF.pdf ITMY-YARM_TF.pdf
Attachment 3: ArmCavNoiseContributions.pdf
ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf
  16126   Fri May 7 11:19:29 2021 Ian MacMillanUpdateCDSSUS simPlant model

I copied c1scx.mdl to the docker to attach to the plant using the commands:

$  ssh nodus.ligo.caltech.edu
[Enter Password]
$  cd opt/rtcds/userapps/release/isc/c1/models/simPlant
$  scp c1scx.mdl controls@c1sim:/home/controls/docker-cymac/userapps
  16125   Thu May 6 16:13:39 2021 AnchalSummaryIMCAngular actuation calibration for IMC mirrors

Here's my first attempt at doing angular actuation calibration for IMC mirrors using the method descibed in /users/OLD/kakeru/oplev_calibration/oplev.pdf by Kakeru Takahashi. The key is to see how much is the cavity mode misaligned from the input mode of beam as the mirrors are moved along PIT or YAW.

There two possible kinds of mismatch:

  • Parallel displacement of cavity mode axis:
    • In this kind of mismatch, the cavity mode is simply away from input mode by some distance \large \beta.
    • This results in transmitted power reduction by the gaussian factor of \large e^{-\frac{\beta^2}{w_0^2}} where \large w_0 is the beam waist of input mode (or nominal waist of cavity).
    • For some mismatch, we can approximate this to
                                                                               \large 1 - \frac{\beta^2}{w_0^2}
  • Angular mismatch of cavity mode axis:
    • The cavity mode axis could be tilted with respect to input mode by some angle \large \alpha.
    • This results in transmitted power reduction by the gaussian factor of \large e^{- \frac{\alpha^2}{\alpha_0^2}}  where \large \alpha_0 is the beam divergence angle of input mode (or nominal waist of cavity) given by \large \frac{\lambda}{\pi w_0}.
    • or some mismatch, we can approximate this to
                                                                                \large 1 - \frac{\alpha^2}{\alpha_0^2}

Kakeru's document goes through cases for linear cavities. For IMC, the mode mismatches are bit different. Here's my take on them:

MC2:

  • MC2 is the easiest case in IMC as it is similar to the end mirror for linear cavity with plane input mirror (the case of which is already studies in sec 0.3.2 in Kaker's document).
  • PIT:
    • When MC2 PIT is changed, the cavity mode simple shifts upwards (or downwards) to the point where the normal from MC2 is horizontal.
    • Since, MC1 and MC3 are plane mirrors, they support this mode just with a different beam spot position, shifted up by \large (R-L)\theta.
    • So the mismatch is simple of the first kind. In my calculations however, I counted the two beams on MC1 and MC3 separately, so the factor is twice as much.
    • Calling the coefficient to square of angular change \large \eta, we get:
                                     \large \eta_{._{2P}} = \frac{2 (R-L)^2}{w_0^2}
    • Here, R is radius of curvature of MC1/3 taken as 21.21m and L is the cavity half-length of IMC taken as 13.545417m.
  • YAW:
    • For YAW, the case is bit more complicated. Similar to PIT, there will be a horizontal shift of the cavity mode by \large (R-L)\theta.
    • But since the MC1 and MC3 mirrors will be fixed, the angle of the two beams from MC1 and MC3 to MC2 will have to shift by \large \theta/2.
    • So the overall coefficient would be:
                                     \large \eta_{._{2Y}} = \frac{2 (R-L)^2}{w_0^2} + \frac{2}{4\alpha_0^2}
    • The factor of 4 in denominator of seconf term on RHS above comes because only half og angular actuation is felt per arm. The factor of 2 in numerator for for the 2 arms.

MC1/3:

  • First, let's establish that the case of MC1 and MC3 is same as the cavity mode must change identically when the two mirrors are moved similarly.
  • YAW:
    • By tilting MC1 by \large \theta, we increase the YAW angle between MC1 and MC3 by \large \theta.
    • Beam spot on both MC1 and MC3 moves by \large (R-L)\theta.
    • The beam angles on both arms get shifted by \large \theta/2.
    • So the overall coefficient would be:
                                     \large \eta_{._{13Y}} = \frac{2 (R-L)^2}{w_0^2} + \frac{2}{4\alpha_0^2}
    • Note, this coefficient is same as MC2, so it si equivalent to moving teh MC2 by same angle in YAW.
  • PIT:
    • I'm not very sure of my caluculation here (hence presented last).
    • Changing PIT on MC1, should change the beam spot on MC2 but not on MC3. Only the angle of MC3-MC2 arm should deflect by \large \theta/2.
    • While on MC1, the beam spot must change by \large (R-L)\theta/2 and the MC1-MC2 arm should deflect by \large \theta/2.
    • So the overall coefficient would be:
                                     \large \eta_{._{13P}} = \frac{(R-L)^2}{4 w_0^2} + \frac{2}{4\alpha_0^2}

Test procedure:

  • We first clicked on MC WFS Relief (on C1:IOO-WFS_MASTER) to reduce the large offsets accumulated on WFS outputs. This script took 10 minutes and reduced the offsets to single digits and IMC remained locked throughout the process.
  • Then we switched off the WFS to freeze the outputs.
  • We moved the MC#_PIT/YAW_OFFSET up and down and measured the C1:IOO-MC_TRANS_SUMFILT_OUT channel as an indicater of IMC mode matching.
  • Attachement 1 are the 6 measurements and there fits to a parabola. Fitting code and plots are thanks to Paco.
  • We got the curvature of parabolas \large \gammafrom these fits in units of 1/cts^2.
  • The \large \eta coefficients calculated above are in units of 1/rad^2.
  • We got the angular actuation calibration from these offsets to physical angular dispalcement in units of rad/cts by \large \sqrt{\gamma / \eta}.
  • AC calibration:
    • I parked the offset to some value to get to the side of parabola. I was trying to reduce transmission from about 14000 cts to 10000-12000 cts in each case.
    • Sent excitation using MC#_ASCPIT/YAW_EXC using awg at 77 Hz and 10000 cts.
    • Measured the cts on transmission channel at 77 Hz. Divided it by 2 and by the dc offset provided. And divided by the amplitude of cts set in excitation. This gives \large \eta_{ac} analogous to above DC case.
    • Then angular actuation calibration at 77 Hz from these offsets to physical angular dispalcement in units of rad/cts by \large \sqrt{\gamma/\eta_{ac}}.
  • Following are the results:
    Optic Act
    Calibration factor at DC [µrad/cts]
    Calibration factor at 77 Hz [prad/cts]
    MC1 PIT 7.931+/-0.029 906.99
    MC1 YAW 5.22+/-0.04 382.42
    MC2 PIT 13.53+/-0.08 869.01
    MC2 YAW 14.41+/-0.21 206.67
    MC3 PIT 10.088+/-0.026 331.83
    MC3 YAW 9.75+/-0.05 838.44

     


  • Note these values are measured with the new settings in effect from 16120. If these are changed, this measurement will not be valid anymore.
  • I believe the small values for MC1 actuation have to do with the fact that coil output gains for MC1 are very weird and small, which limit the actuation strength.
  • TAbove the resonance frequencies, they will fall off by 1/f^2 from the DC value. I've confirmed that the above numbers are of correct order of magnitude atleast.
  • Please let me know if you can point out any mistakes in the calculations above.
Attachment 1: IMC_Ang_Act_Cal_Kakeru_Tests.pdf
IMC_Ang_Act_Cal_Kakeru_Tests.pdf IMC_Ang_Act_Cal_Kakeru_Tests.pdf IMC_Ang_Act_Cal_Kakeru_Tests.pdf IMC_Ang_Act_Cal_Kakeru_Tests.pdf IMC_Ang_Act_Cal_Kakeru_Tests.pdf IMC_Ang_Act_Cal_Kakeru_Tests.pdf
  16124   Thu May 6 16:13:24 2021 Ian MacMillanUpdateCDSSUS simPlant model

When using mdl2adl I was getting the error:

$  cd /home/controls/mdl2adl
$  ./mdl2adl x1sup.mdl
error: set $site and $ifo environment variables

to set these in the terminal use the following commands:

$  export site=tst
$  export ifo=x1

On most of the systems, there is a script that automatically runs when a terminal is opened that sets these but that hasn't been added here so you must run these commands every time you open the terminal when you are using mdl2adl.

  16122   Wed May 5 15:11:54 2021 Ian MacMillanUpdateCDSSUS simPlant model

I added the IPC parts back to the plant model so that should be done now. It looks like this again here.

I can't seem to find the control model which should look like this. When I open sus_single_control.mdl, it just shows the C1_SUS_SINGLE_PLANT.mdl model. Which should not be the case.

  16121   Wed May 5 13:05:07 2021 ChubUpdateGeneralchassis delivery from De Leone

Assembled chassis from De Leone placed in the 40 Meter Lab, along the west wall and under the display pedestal table.  The leftover parts are in smaller Really Useful boxes, also on the parts pile along the west wall.

Attachment 1: de_leone_del_5-5-21.jpg
de_leone_del_5-5-21.jpg
  16120   Wed May 5 09:04:47 2021 AnchalUpdateSUSNew IMC Suspension Damping Gains uploaded for long term testing

We have uploaded the new damping gains on all the suspensions of IMC. This completes changing all the configuration to as mentioned in 16066 and 16072. The old setting can be restored by running python3 /users/anchal/20210505_IMC_Tuned_SUS_with_Gains/restoreOldConfigIMC.py from allegra or donatella.

GPSTIME: 1304265872

UTC May 05, 2021 16:04:14 UTC
Central May 05, 2021 11:04:14 CDT
Pacific May 05, 2021 09:04:14 PDT

 

  16119   Tue May 4 19:14:43 2021 YehonathanUpdateGeneralOSEMs from KAGRA

I put the box containing the untested OSEMs from KAGRA near the south flow bench on the floor.

  16118   Tue May 4 14:55:38 2021 Ian MacMillanUpdateCDSSUS simPlant model

After a helpful meeting with Jon, we realized that I have somehow corrupted the sitemap file. So I am going to use the code Chris wrote to regenerate it. 

Also, I am going to connect the controller using the IPC parts. The error that I was having before had to do with the IPC parts not being connected properly. 

  16117   Tue May 4 11:43:09 2021 Anchal, PacoUpdateLSCIMC WFS noise contribution in arm cavity length noise

We redid the WFS noise injection test and have compiled some results on noise contribution in arm cavity noise and IMC frequency noise due to angular noise of IMC.


Attachment 1: Shows the calibrated noise contribution from MC1 ASCPIT OUT to ARM cavity length noise and IMC frequency noise.

  • For calibrating the cavity length noise signals, we sent 100 cts 100Hz sine excitation to ITMX/Y_LSC_EXC, used actuator calibration for them as 2.44 nm/cts from 13984, and measured the peak at 100 hz in time series data. We got calibration factors: ETMX-LSC_OUT: 60.93 pm/cts , and ETMY-LSC_OUT: 205.0 pm/cts.
  • For converting IMC frequency noise to length noise, we used conversion factor given by \lambda L / c where L is 37.79m and lambda is wavelength of light.
  • For converting MC1 ASCPIT OUT cts data to frequency noise contributed to IMC, we sent 100,000 amplitude bandlimited noise (see attachment 3 for awggui config) from 25 Hz to 30 Hz at C1:IOO-MC1_PIT_EXC. This noise was seen at both MC_F and ETMX/Y_LSC_OUT channels. We used the noise level at 29 Hz to get a calibration for MC1_ASCPIT_OUT to IMC Frequency in Hz/cts. See Attachment 2 for the diaggui plots.
  • Once we got the calibration above, we measured MC1_ASCPIT_OUT power spectrum without any excitaiton and multiplied it with the calibration factor.
  • However, something must be wrong because the MC_F noise in length units is coming to be higher than cavity length noise in most of the frequency band.
    • It can be due to the fact that control signal power spectrum is not exactly cavity length noise at all frequencies.  That should be only above the UGF of the control loop (we plan to measure that in afternoon).
    • Our calibration for ETMX/Y_LSC_OUT might be wrong.
Attachment 1: ArmCavNoiseContributions.pdf
ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf ArmCavNoiseContributions.pdf
Attachment 2: IOO-MC1_PIT_NoiseInjTest2.pdf
IOO-MC1_PIT_NoiseInjTest2.pdf IOO-MC1_PIT_NoiseInjTest2.pdf
Attachment 3: IOO-MC1_PIT_NoiseInjTest_AWGGUI_Config.png
IOO-MC1_PIT_NoiseInjTest_AWGGUI_Config.png
  16116   Tue May 4 07:38:36 2021 JonUpdateCDSI/O Chassis Assembly

IOP models created

With all the PCIe issues now resolved, yesterday I proceeded to build an IOP model for each of new FEs. I assigned them names and DCUIDs consist with the 40m convention, listed below. These models currently exist on only the cloned copy of /opt/rtcds running on the test stand. They will be copied to the main network disk later, once the new systems are fully tested.

Model Host CPU DCUID
c1x06 c1bhd 1 23
c1x07 c1sus2 1 24

The models compile and install successfully. The RCG runtime diagnostics indicate that all is working except for the timing synchronization and DAQD data transmission. This is as expected because neither of these have been set up yet.

Timing system set-up

The next step is to provide the 65 kHz clock signals from the timing fanout via LC optical fiber. I overlooked the fact that an SPX optical transceiver is required to interface the fiber to the timing slave board. These were not provided with the timing slaves we received. The timing slaves require a particular type of transceiver, 100base-FX/OC-3, which we did not have on hand. (For future reference, there is a handy list of compatible transceivers in E080541, p. 14.) I placed a Digikey order for two Finisar FTLF1217P2BTL, which should arrive within two days.

Attachment 1: Screen_Shot_2021-05-03_at_4.16.06_PM.png
Screen_Shot_2021-05-03_at_4.16.06_PM.png
  16115   Mon May 3 23:28:56 2021 KojiSummaryGeneralWeird gas leakagr kind of noise in 40m control room

I also noticed some sound in the control room. (didn't open the MP3 yet)

I'm afraid that the hard disk in the control room iMac is dying.

 

  16114   Mon May 3 20:36:46 2021 Yehonathan, JonUpdateCDSUpdated c1auxey wiring plan

It seemed like the BIO channels were not working, both the inputs and the outputs. The inputs were working on the windows machine though. That is, when we shorted the BIO channel to the return, or put 0V on it, we could see the LED turn on on the I/O testing screen and when we ramped up the voltage above 3 the LED turned off. This is the expected behavior from a sinking digital input. However, the EPICs caget didn't show any change. All the channels were stuck on Disabled.

We checked the digital outputs by connecting the channels to a fluke. Initially, the fluke showed 13V. We tried to toggle the digital output channels with caput and that didn't work. We checked the outputs with the windows software. For that, we needed to stop the Modbus. To our surprise, the windows software was not able to flip the channels either. We realized that this BIO Acromag unit is probably defective. We replaced it with a different unit and put a warning sticker on the defective unit. Now, the digital outputs were working as expected. When we turned them on the voltage output dropped to 0V. We checked the channels with the EPICs software. We realized that these channels were locked with the closed loop definition. We turned on the channels tied to these output channels (watchdog and toggles) and it worked. The output channels can be flipped with the EPICs software. We checked all the digital output channels and fixed some wiring issues along the way.

The digital input channels were still not working. This is a software issue that we will have to deal with later.

(Yehonathan) Rana noticed that the BNC leads on the chassis front panel didn't have isolation on them so I redid them with shrinking tubes.

  16113   Mon May 3 18:59:58 2021 AnchalSummaryGeneralWeird gas leakagr kind of noise in 40m control room

For past few days, a weird sound of decaying gas leakage comes in the 40m control room from the south west corner of ceiling. Attached is an audio capture. This comes about every 10 min or so. 

Attachment 1: 40mNoiseFinal.mp3
  16112   Mon May 3 17:28:58 2021 Anchal, Paco, RanaUpdateLSCIMC WFS noise contribution in arm cavity length noise

Rana came and helped us figure us where to inject the noise. Following are the characteristics of the test we did:

  • Inject normal noise at C1:IOO-MC1_PIT_EXC using AWGGUI.
  • Excitation amplitude of 54321 in band 12-37Hz with Cheby1 8th order bandpass filter with same limits.
  • Look at power spectrum of C1:IOO-MC_F_DQ, C1:IOO-WFS1-PIT_OUT_DQ and the C1:IOO-MC1_PIT_EXC itself.
  • Increased the gain of the noise excitation until we see some effect in MC_F.
  • Diaggui also showed coherence plot in the bottom, which let's us have an estimate of how much we need to go further.

Attachment 1 shows a screenshot with awggui and diaggui screens displaying the signal in both angular and longitudinal channels.

Attachment 2 shows the analogous screenshot for MC2.

 

Attachment 1: excitationoftheMCanglessothatwecanseesomethingdotpng.png
excitationoftheMCanglessothatwecanseesomethingdotpng.png
Attachment 2: excitationoftheMCanglessothatwecanseesomethingdotpngbutthistimeitsMC2.png
excitationoftheMCanglessothatwecanseesomethingdotpngbutthistimeitsMC2.png
  16111   Mon May 3 16:49:04 2021 YehonathanUpdateBHDSOS assembly

I found a "vice" in the cleanroom (attachment 1). I used it to push dowel pins into the last suspension block using some alcohol as a lubricant.

I then assembled the 7th and last suspension tower (attachment 2).

Things that need to be done:

1. Push Viton tips into vented screws and assemble the earthquake stops.

2. Glue magnets to dumbells.

Attachment 1: 20210503_161422.png
20210503_161422.png
Attachment 2: 20210503_161456.jpg
20210503_161456.jpg
  16110   Mon May 3 16:24:14 2021 AnchalUpdateSUSIMC Suspension Damping Gains Test Repeated with IMC unlocked

We repeated the same test with IMC unlocked. We had found these gains when IMC was unlocked and their characterization needs to be done with no light in the cavity. attached are the results. Everything else is same as before.

Quote:

With the input matrix, coil ouput gains and F2A filters loaded as in 16091, I tested the suspension loops' step response to offsets in LSC, ASCPIT and ASCYAW channels, before and after applying the "new damping gains" mentioned in 16066 and 16072. If these look better, we should upload the new (higher) damping gains as well. This was not done in 16091.


Note that in the plots, I have added offsets in the different channels to plot them together, hence the units are "au".


Edit Tue May 4 14:43:48 2021 :

  • Adding zoomed in plots to show first 25s after the step.
  • MC1:
    • Our improvements by new gains are only modest.
    • This optic needs a more careful coil balancing first.
    • Still the ring time is reduced to about 5s for all step responses as opposed to 10s at old gains.
  • MC2:
    • The first page of MC2 might be bit misleading. We have not changed the damping gain for SUSPOS channel, so the longer ringing is probably just an artifact of somthing else. We didn't retake data.
    • In PIT and YAW where we increased the gain by a factor of 3, we see a reduction in ringing lifetime by about half.
  • MC3:
    • We saw the most optimistic improvement on this optic.
    • The gains were unusually low in this optic, not sure why.
    • By increasing SUSPOS gain from 200 to 500, we saw a reduction of ringing halftime from 7-8s to about 2s. Improvements are seen in other DOFs as well.
    • You can notice rightaway that YAW of MC3 keeps oscillating near resonance (about 1 Hz). Maybe more careful feedback shaping is required here.
    • In SUSPIT, we increased gain from 12 to 35 and saw a good reduction in both ringing time and initial amplitude of ringing.
    • In SUSYAW, we only increased the gain to 12 from 8, which still helped a lot in reducing big ringing step response to below 5s from about 12s.

Overall, I would recommend setting the new gains in the suspension loops as well to observe long term effects too.

Attachment 1: MC1_SusDampGainTest.pdf
MC1_SusDampGainTest.pdf MC1_SusDampGainTest.pdf MC1_SusDampGainTest.pdf
Attachment 2: MC2_SusDampGainTest.pdf
MC2_SusDampGainTest.pdf MC2_SusDampGainTest.pdf MC2_SusDampGainTest.pdf
Attachment 3: MC3_SusDampGainTest.pdf
MC3_SusDampGainTest.pdf MC3_SusDampGainTest.pdf MC3_SusDampGainTest.pdf
  16109   Mon May 3 13:35:12 2021 Ian MacMillanUpdateCDSSUS simPlant model

When the cymac is started it gives me a list of channels shown below. 

 $  Initialized TP interface node=8, host=98e93ecffcca
 $  Creating X1:DAQ-DC0_X1IOP_STATUS
 $  Creating X1:DAQ-DC0_X1IOP_CRC_CPS
 $  Creating X1:DAQ-DC0_X1IOP_CRC_SUM
 $  Creating X1:DAQ-DC0_X1SUP_STATUS
 $  Creating X1:DAQ-DC0_X1SUP_CRC_CPS
 $  Creating X1:DAQ-DC0_X1SUP_CRC_SUM

But when I enter it into the Diaggui I get an error:

The following channel could not be found:
X1:DAQ-DC0_X1SUP_CRC_CPS

My guess is that need to connect to the Diaggui to something that can access those channels. I also need to figure out what those channels are.

  16108   Mon May 3 09:14:01 2021 Anchal, PacoUpdateLSCIMC WFS noise contribution in arm cavity length noise

Lock ARMs

  • Try IFO Configure ! Restore Y Arm (POY) and saw XARM lock, not YARM. Looks like YARM biases on ITMY and ETMY are not optimal, so we slide C1:SUS-ETMY_OFF from 3.0 --> -14.0 and watch Y catch its lock.
  • Run ASS scripts for both arms and get TRY/TRX ~ 0.95
    • We ran X, then Y and noted that TRX dropped to ~0.8 so we ran it again and it was well after that. From now on, we will do Y, then X.

WFS1 noise injection

  • Turn WFS limits off by running switchOffWFSlims.sh
  • Inject broadband noise (80-90 Hz band) of varying amplitudes from 100 - 100000 counts on C1:IOO-WFS1_PIT_EXC
  • After this we try to track its propagation through various channels, starting with
    • C1:LSC-XARM_IN1_DQ / C1:LSC-YARM_IN1_DQ
    • C1:SUS-ETMX_LSC_OUT_DQ / C1:SUS-ETMY_LSC_OUT_DQ
    • C1:IOO-MC_F_DQ
    • C1:SUS-MC1_**COIL_OUT / C1:SUS-MC2_**COIL_OUT / C1:SUS-MC3_**COIL_OUT
    • C1:IOO-WFS1_PIT_ERR / C1:IOO-WFS1_YAW_ERR
    • C1:IOO-WFS1_PIT_IN2

** denotes [UL, UR, LL, LR]; the output coils.

  • Attachment 1 shows the power spectra with IMC unlocked
  • Attachment 2 shows the power spectra with the ARMs (and IMC) locked
Attachment 1: WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdf
WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdf WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdf WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdf WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdf WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdf WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdf WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdf WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdf WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdf
Attachment 2: WFS1_PIT_Noise_Inj_Test_ARM_locked.pdf
WFS1_PIT_Noise_Inj_Test_ARM_locked.pdf WFS1_PIT_Noise_Inj_Test_ARM_locked.pdf WFS1_PIT_Noise_Inj_Test_ARM_locked.pdf WFS1_PIT_Noise_Inj_Test_ARM_locked.pdf WFS1_PIT_Noise_Inj_Test_ARM_locked.pdf WFS1_PIT_Noise_Inj_Test_ARM_locked.pdf WFS1_PIT_Noise_Inj_Test_ARM_locked.pdf WFS1_PIT_Noise_Inj_Test_ARM_locked.pdf WFS1_PIT_Noise_Inj_Test_ARM_locked.pdf
  16107   Fri Apr 30 19:18:51 2021 Yehonathan, JonUpdateCDSUpdated c1auxey wiring plan

We finished the installation procedure on the c1auxey1 host machine. There were some adjustments that had to be made for Debian 10. The slow machine wiki page has been updated.

A test database file was made were all the channel names were changed from C1 to C2 in order to not interfere with the existing channels.

We starting testing the channels one by one to check the wiring and the EPICs software. We found some misswirings and fixed them.

Channel Name Type EPICs Test Acromag windows software test
C2:SUS-ETMY_ULPDMon AI Pass Pass
C2:SUS-ETMY_URPDMon AI Pass Pass
C2:SUS-ETMY_LLPDMon AI Pass Pass
C2:SUS-ETMY_SPDMon AI Pass Pass
C2:SUS-ETMY_LRPDMon AI Pass Pass
C2:SUS-ETMY_ULVMon AI Pass Pass
C2:SUS-ETMY_URVMon AI Pass Pass
C2:SUS-ETMY_LLVMon AI Pass Pass
C2:SUS-ETMY_SideVMon AI Pass Pass
C2:SUS-ETMY_LRVMon AI Pass Pass

Its getting late. I'll continue with the rest of the channels on Monday.

Notice that for all the AI channels the RTN was disconnected while testing.

 

 

 

 

 

 

  16106   Fri Apr 30 12:52:14 2021 Ian MacMillanUpdateCDSSUS simPlant model

Now that the model is finally compiled I need to make an medm screen for it and put it in the c1sim:/home/controls/docker-cymac/userapps/medm/ directory. 

But before doing that I really want to test it using the autogenerated medm screens which are in the virtual cymac in the folder /opt/rtcds/tst/x1/medm/x1sup. In Jon's post he said that I can use the virtual path for sitemap after running $ eval $(./env_cymac)

  16105   Fri Apr 30 00:20:30 2021 gautamUpdateCDSF2A Filters double check

The SDF system is supposed to help with restoring the correct settings, complementary to burt. My personal opinion is that there is no need to commit these filters to SDF until we're convinced that they help with the locking / noise performance.

Quote:

I double checked today and the F2A filters in the output matrices of MC1, MC2 and MC3 in the POS column are ON. I do not get what SDF means? Did we need to add these filters elsewhere

  16104   Fri Apr 30 00:18:40 2021 gautamSummaryLSCStart of measuring IMC WFS noise contribution in arm cavity length noise

This is the actuator calibration. For the error point calibration, you have to look at the filter in the calibration model. I think it's something like 8e-13m/ct for POX and similar for POY.

Quote:

I calibrated the control arms signals by 2.44 nm/cts calibration factor directly picked up from 13984.

  16103   Thu Apr 29 19:55:45 2021 YehonathanUpdateCDSUpdated c1auxey wiring plan

We received a stock of DB9 male feed-through connectors. That allowed me to complete the remaining wiring on the c1auxey Acromag chassis. The only thing left to be done is the splicing to the RTS.

 

  16102   Thu Apr 29 18:53:33 2021 AnchalUpdateSUSIMC Suspension Damping Gains Test

With the input matrix, coil ouput gains and F2A filters loaded as in 16091, I tested the suspension loops' step response to offsets in LSC, ASCPIT and ASCYAW channels, before and after applying the "new damping gains" mentioned in 16066 and 16072. If these look better, we should upload the new (higher) damping gains as well. This was not done in 16091.


Note that in the plots, I have added offsets in the different channels to plot them together, hence the units are "au".

Attachment 1: MC1_SUSDampGainTest.pdf
MC1_SUSDampGainTest.pdf MC1_SUSDampGainTest.pdf MC1_SUSDampGainTest.pdf
Attachment 2: MC2_SUSDampGainTest.pdf
MC2_SUSDampGainTest.pdf MC2_SUSDampGainTest.pdf MC2_SUSDampGainTest.pdf
Attachment 3: MC3_SUSDampGainTest.pdf
MC3_SUSDampGainTest.pdf MC3_SUSDampGainTest.pdf MC3_SUSDampGainTest.pdf
  16101   Thu Apr 29 17:51:19 2021 AnchalSummaryLSCStart of measuring IMC WFS noise contribution in arm cavity length noise

t Both arms were locked simply by using IFO > Configure > ! (YARM) > Restore YARM. I had to use ASS to improve the TRX/TRY to ~0.95.

I measured C1:LSC-XARM_IN1_DQ and C1:LSC-YARM_IN1_DQ while injecting band limited noise in C1:IOO-WFS1_PIT_EXC using uniform noise with amplitude 1000 along with filter defined by string: cheby1("BandPass",4,1,80,100). I calibrated the control arms signals by 2.44 nm/cts calibration factor directly picked up from 13984.

For the duration of this test, all LIMIT switches in the WFS loops were switched OFF.

I do not see any affect on the arm control signal power spectrums with or without the noise injection. Attachement 1 shows the PSD along with PSD of the injection site IN2 signal. I must be doing something wrong, so would like feedback before I go further.

Attachment 1: WFS1_PIT_exc_80-100Hz_Arms_ASD.pdf
WFS1_PIT_exc_80-100Hz_Arms_ASD.pdf
  16100   Thu Apr 29 17:43:48 2021 AnchalUpdateCDSF2A Filters double check

I double checked today and the F2A filters in the output matrices of MC1, MC2 and MC3 in the POS column are ON. I do not get what SDF means? Did we need to add these filters elsewhere?

Quote:

The IMC suspension team should double check their filters are on again. I am not familiar with the settings and I don't think they've been added to the SDF.

Attachment 1: F2AFiltersON.png
F2AFiltersON.png
  16099   Thu Apr 29 17:43:16 2021 KojiUpdateCDSRFM

The other day I felt hot at the X end. I wondered if the Xend A/C was off, but the switch right next to the SP table was ON (green light).
I could not confirm if the A/C was actually blowing or not.

  16098   Thu Apr 29 16:35:51 2021 YehonathanUpdateCDSUpdated c1auxey wiring plan

I installed the EPICs base, asyn and modbus modules according to Jon's instructions.

Since the modbus configurations files were already writtten for c1auxey1 (see elog 15292) the only thing I did was to change the IP addresses in ETMYaux.cmd to match the actual assigned IPs.

I followed the rest of the instructions as written.

The modbus service was activated succesfully.

The only thing left to do is to change ETMYaux.db to reflect to new channels that were added. I believe these are BI channels named C1:SUS-ETMY_xx_ENABLEMon.

 

  16097   Thu Apr 29 15:11:33 2021 gautamUpdateCDSRFM

The problem here was that the RFM errors cropped up again - seems like it started ~4am today morning judging by TRX trends. Of course without the triggering signal the arm cavity couldn't lock. I rebooted everything (since just restarting the rfm senders/receivers did not do the trick), now arm locking works fine again. It's a bit disappointing that the Rogue Master setting did not eliminate this problem completely, but oh well...

It's kind of cool that in this trend view of the TRX signal, you can see the drift of the ETMX suspension. The days are getting hot again and the temp at EX can fluctuate by >12C between day and night (so the "air-conditioning" doesn't condition that much I guess 😂 ), and I think that's what drives the drift (idk what the transfer function to the inside of the vacuum chamber is but such a large swing isn't great in any case). Not plotted here but i hypothesize TRY levels will be more constant over the day (modulo TT drift which affects both arms).

The IMC suspension team should double check their filters are on again. I am not familiar with the settings and I don't think they've been added to the SDF.

Attachment 1: RFM_errs.png
RFM_errs.png
Attachment 2: Screenshot_2021-04-29_15-12-56.png
Screenshot_2021-04-29_15-12-56.png
  16096   Thu Apr 29 13:41:40 2021 Ian MacMillanUpdateCDSSUS simPlant model

To add the required library: put the .mdl file that contains the library into the userapps/lib folder. That will allow it to compile correctly

I got these errors:

Module ‘mbuf’ symvers file could not be found.
Module ‘gpstime’ symvers file could not be found.
***ERROR: IPCx parameter file /opt/rtcds/zzz/c1/chans/ipc/c1.ipc not found
make[1]: *** [Makefile:30: c1sup] Error 2
make: *** [Makefile:35: c1sup] Error 1

I removed all IPC parts (as seen in Attachment 1) and that did the trick. IPC parts (Inter-Process Communication) were how this model was linked to the controller so I don't know how exactly how I can link them now. 

I also went through the model and grounded all un-attached inputs and outputs. Now the model compiles

Also, The computer seems to be running very slowly in the past 24 hours. I know Jon was working on it so I'm wondering if that had any impact. I think it has to do with the connection speed because I am connected through X2goclient. And one thing that has probably been said before but I want to note again is that you don't need a campus VPN to access the docker.

Attachment 1: Non-IPC_Plant.pdf
Non-IPC_Plant.pdf
  16095   Thu Apr 29 11:51:27 2021 AnchalSummaryLSCStart of measuring IMC WFS noise contribution in ar cavity length noise

Tried locking the arms

  • Ran IFO > Configure > ! (YARM) > Restore YARM. Nothing happened.
  • Trying to align through tip-tilt:
    • Previous values: TT1: PIT: -1.7845, YAW: -0.2775; TT2: PIT: -1.3376, YAW: -0.0436
    • Couldn't get flashing of light in the arms at all.
    • Restored values to previous values.
  • Noticed that ITMY OPLEV YAWW Error has gone very high overnight while other oplevs remained the same.
  • Trying to change the C1:SUS-ITMY_YAW_OFFSET to bring this oplev yaw error back to near zero.
  • Changed C1:SUS-ITMY_YAW_OFFSET from -34 to 50. OPLEV_YEROR reduced to around -23 from -70.
  • Same thing with BS PIT. OPLEV_PERROR is highlighted in red at -52.
  • Changing C1:SUS-BS_PIT_OFFSET from 55 to 30. This brought OPLEV_PERROR to -15 from -52.
  • Trying to align PRM by changing C1:SUS-PRM_PIT_OFFSET and C1:SUS-PRM_YAW_OFFSET.
  • Inital values: C1:SUS-PRM_PIT_OFFSET: -20 , C1:SUS-PRM_YAW_OFFSET: 39.

Did the WFS step response test on IMC in between while waiting for help. See 16094.


Back to trying arm locking

  • Tried IFO > Configure > ! (XARM) > Restore YARM. Nothing happened.
  • Tried IFO > Configure > ! (YARM) > Restore YARM. Nothing happened again.
  • Tried Movie Capture of AS screen from VIDEO > Movie Capture > AS. But the script failed due to module not present error.

PMC got unlocked

  • Infront of me, PMC got unlocked. I did not go to PMC locking the screen at all since morning.
  • I opened the C1PSL_PMC screen. The PSL Autolocker blinker is not blinking but the switch is set to Enable. 
  • I do not see any automatic effort happening for regaining lock at PMC.
  • I'll try it manually. I was able to get the PMC locked again. C1:PSL-PMC_PMCTRANSPD is showing 0.761 V and C1:PSL-PMC_RFPDDC is showing 0.053 V.
  • Now IMC auto locker seems to be trying to get the lock acquired.
  • It acquired a lock a few times but struggled to keep it on. I reduced C1:IOO-WFS_GAIN to 0 and then the lock could stay on. Seemed like the accumulated offsets were not good.
  • So I cleared the history on WFS1, TRANS and WFS2 filter banks and then ramped the WFS overall gain (C1:IOO-WFS_GAIN) back to 1 and now IMC seems to stay locked in a stable configuration.
  • However, I still don't know what caused the PMC to get unlocked in the first place. Did my repeated arm locking attempts did something to the main laser frequency?

Back to trying arm locking

  • Tried IFO > Configure > ! (YARM) > Restore YARM again. Nothing happened again.
  16094   Thu Apr 29 10:52:56 2021 AnchalUpdateSUSIMC Trans QPD and WFS loops step response test

In 16087 we mentioned that we were unable to do a step response test for WFS loop to get an estimate of their UGF. The primary issue there was that we were not putting the step at the right place. It should go into the actuator directly, in this case, on C1:SUS-MC2_PIT_COMM and C1:SUS-MC2_YAW_COMM. These channels directly set an offset in the control loop and we can see how the error signals first jump up and then decay back to zero. The 'half-time' of this decay would be the inverse of the estimated UGF of the loop. For this test, the overall WFS loops gain,  C1:IOO-WFS_GAIN was set to full value 1. This test is performed in the changed settings uploaded in 16091.

I did this test twice, once giving a step in PIT and once in YAW.

Attachment 1 is the striptool screenshot for when PIT was given a step up and then step down by 0.01.

  • Here we can see that the half-time is roughly 10s for TRANS_PIT and WFS1_PIT corresponding to roughly 0.1 Hz UGF.
  • Note that WFS2 channels were not disturbed significantly.
  • You can also notice that third most significant disturbance was to TRANS_YAW actually followed by WF1 YAW.

Attachment 2 is the striptool screenshot when YAW was given a step up and down by 0.01. Note the difference in x-scale in this plot.

  • Here, TRANS YAW got there greatest hit and it took it around 2 minutes to decay to half value. This gives UGF estimate of about 10 mHz!
  • Then, weirdly, TRANS PIT first went slowly up for about a minutes and then slowly came dome in a half time of 2 minutes again. Why was PIT signal so much disturbed by the YAW offset in the first place?
  • Next, WFS1 YAW can be seen decaying relatively fast with half-life of about 20s or so.
  • Nothing else was disturbed much.

  • So maybe we never needed to reduce WFS gain in our measurement in 16089 as the UGF everywhere were already very low.
  • What other interesting things can we infer from this?
  • Should I sometime repeat this test with steps given to MC1 or MC3 optics?
Attachment 1: PIT_OFFSET_ON_MC2.png
PIT_OFFSET_ON_MC2.png
Attachment 2: YAW_STEP_ON_MC2_complete.png
YAW_STEP_ON_MC2_complete.png
  16093   Thu Apr 29 10:51:35 2021 JonUpdateCDSI/O Chassis Assembly

Summary

Yesterday I unpacked and installed the three 18-bit DAC cards received from Hanford. I then repeated the low-level PCIe testing outlined in T1900700, which is expanded upon below. I did not make it to DAC-ADC loopback testing because these tests in fact revealed a problem with the new hardware. After a combinatorial investigation that involved swapping cards around between known-to-be-working PCIe slots, I determined that one of the three 18-bit DAC cards is bad. Although its "voltage present" LED illuminates, the card is not detected by the host in either I/O chassis.

I installed one of the two working DACs in the c1bhd chassis. This now 100% completes this system. I installed the other DAC in the c1sus2 chassis, which still requires four more 18-bit DACs. Lastly, I reran the PCIe tests for the final configurations of both chassis.

PCIe Card Detection Tests

For future reference, below is the set of command line tests to verify proper detection and initialization of ADC/DAC/BIO cards in I/O chassis. This summarizes the procedure described in T1900700 and also adds the tests for 18-bit DAC and 32-channel BO cards, which are not included in the original document.

Each command should be executed on the host machine with the I/O chassis powered on:

$ sudo lspci -v | grep -B 1 xxxx

where xxxx is a four-digit device code given in the following table.

Device Device Code
General Standards 16-bit ADC 3101
General Standards 16-bit DAC 3120
General Standards 18-bit DAC 3357
Contec 16-channel BIO 8632
Contec 32-channel BO 86e2
Dolphin IPC host adapter 0101

The command will return a two-line entry for each PCIe device of the specified type that is detected. For example, on a system with a single ADC this command should return:

10:04.0 Bridge: PLX Technology, Inc. PCI9056 32-bit 66MHz PCI IOBus Bridge (rev ac)
             Subsystem: PLX Technology, Inc. Device 3101
  16092   Wed Apr 28 18:56:57 2021 Yehonathan, JonUpdateCDSUpdated c1auxey wiring plan

We took a Supermicro from the lab (along with a keyboard, a mouse, and a screen taken from a table on the Y arm) and placed it near the Acromag chassis.

We installed Debian 10 on the machine. I followed the steps on the slow machine wiki for setting up the host machine. Some steps had to be updated. Most importantly, in the new Debian, the network interfaces are given random names like enp3s0 and enp4s0 instead of eth0 and eth1. I updated the wiki accordingly.

To operate the chassis using one 15V source I disconnected the +24V cable from the Acromag units and jumpered the +15V wire into the power input instead. I started up the Acromags. They draw 0.7A. I connected an Ethernet cable to the front interface. I checked that all the Acromags are connected to the local network of the host machine by pinging them one by one.

 

  16091   Wed Apr 28 17:09:11 2021 AnchalUpdateSUSTuned Suspension Parameters uploaded for long term behavior monitoring

I have uploaded all the new settings mentioned in 16066 and 16072. The settings were uploaded through a single script present at anchal/20210428_IMC_Tuned_Suspension/uploadNewConfigIMC.py. The settings can be reverted back to old settings through anchal/20210428_IMC_Tuned_Suspension/restoreOldConfigIMC.py. Both these scripts can be run only through python3 in donatella or allegra.


GPSTIME of new settings: 1303690144


New settings include:

  • New input matrices for MC1 and MC2.
  • New Output coil gains for AC balancing on all three optics.
  • Switching ON the FM8 filter modulae (Q=3 F2A filter) in POS column on output matrix of all optics.

We'll wait and watch the performance through summary pages and check back the performance on Monday.

  16090   Wed Apr 28 11:31:40 2021 JonUpdateVACEmpty N2 Tanks

I checked out what happened on c1vac. There are actually two independent monitoring codes running:

  1. The interlock service, which monitors the line directly connected to the valves.
  2. A seaparate convenience mailer, running as a cronjob, that monitors the tanks.

The interlocks did not trip because the low-pressure delivery line, downstream of the dual-tank regulator, never fell below the minimum pressure to operate the valves (65 PSI). This would have eventually occurred, had Jordan been slower to replace the tanks. So I see no problem with the interlocks.

On the other hand, the N2 mailer should have sent an email at 2021-04-18 15:00, which was the first time C1:Vac-N2T1_pressure dropped below the 600 PSI threshold. N2check.log shows these pressures were recorded at this time, but does not log that an email was sent. Why did this fail? Not sure, but I found two problems which I did fix:

  • One was that the code was checking the sensor on the low-pressure side (C1:Vac-N2_pressure; nominally 75 PSI) against the same 600 PSI threshold as the tanks. This channel should either be removed or a separate threshold (65 PSI) defined just for it. I just removed it from the list because monitoring of this channel is redundant with the interlock service. This does not explain the failure to send an email.
  • The second issue was that the pyN2check.sh script appeared to be calling Python 3 to run a Python 2 script. At least that was the situation when I tested it, and this was causing it to fail partway through. This might well explain the problem with no email. I explicitly set python --> python2 in the shell script.

The code then ran fine for me when I retested it. I don't see any further issues.

Quote:

Installed T2 today, and leaked checked the entire line. No issues found. It could have been a bad valve on the tank itself. Monitored T2 pressure for ~2 hours to see if there was any change. All seems ok.

Quote:

When I came into the lab this morning, I noticed that both N2 tanks were empty. I had swapped one on Friday (4-16-21) before I left the lab. Looking at the logs, the right tank (T2) sprung a leak shortly shortly after install. I leak checked the tank coupling after install but did not see a leak. There could a leak further down the line, possibly at the pressure transducer.

The left tank (T1) emptied normally over the weekend, and I quickly swapped the left tank for a full one, and is curently at ~2700 psi. It was my understanding that if both tanks emptied, V1 would close automatically and a mailer would be sent out to the 40m group. I did not receive an email over the weekend, and I checked the Vac status just now and V1 was still open.

I will keep an eye on the tank pressure throughout the day, and will try to leak check the T2 line this afternoon, but someone should check the vacuum interlocks and verify.

 

  16089   Wed Apr 28 10:56:10 2021 Anchal, PacoUpdateSUSIMC Filters diagnosed

Good morning!

We ran the f2a filter test for MC1, MC2, and MC3.


Filters

The new filters differ from previous versions by a adding non-unity Q factor for the pole pairs as well.

\frac{f^2 - i \frac{f_z}{Q}f + f_z^2}{f^2 - i \frac{f_0}{Q}f + f_0^2}
This in terms of zpk is: [ [zr + i zi, zr - i zi], [pr + i pi, pr - i pi], 1] where
z_r = -\frac{f_z}{2Q}, \quad z_i = f_z \sqrt{1 - \frac{1}{4Q^2}}, \quad p_r = -\frac{f_0}{2Q}, \quad p_i = f_0 \sqrt{1 - \frac{1}{4Q^2}}, \quad f_z = f_0 \sqrt{G_{DC}}

  • Attachment #1 shows the filters for MC1 evaluated for Q=3, 7,and 10.
  • Attachment #2 shows the filters for MC2 evaluated for Q=3, 7, and 10.
  • Attachment #3 shows the filters for MC3 evaluated for Q=3, 7, and 10.
  • Attachment #4 shows the bode plots generated by foton after uploading for Q=3 case.

We uploaded all these filters using foton, into the three last FM slots on the POS output gain coil.


Tests

We ran tests on all suspended optics using the following (nominal) procedure:

  1. Upload new input matrix
  2. Lower the C1:IOO-WFS_GAIN to 0.05.
  3. Upload AC coil balancing gains.
  4. Take ASD for the following channels:
    • C1:IOO-MC_TRANS_PIT_IN1
    • C1:IOO-MC_TRANS_YAW_IN1
    • C1:IOO-MC_WFS1_PIT_IN1
    • C1:IOO-MC_WFS1_YAW_IN1
    • C1:IOO-MC_WFS2_PIT_IN1
    • C1:IOO-MC_WFS2_YAW_IN1
  5. For the following combinations:
    • No excitation** + no filter
    • No excitation + filter
    • Excitation + no filter
    • Excitation + filter

** Excitation = 0.05 - 3.5 Hz uniform noise, 100 amplitude, 100 gain


Plots

  • Attachment 5-7 give the test results for MC1, MC2 and MC3.
  • In each pdf, the three pages show ASD of TRANS QPD, WFS1 and WFS2 channels' PIT and YAW, respectively.
  • Red/blue correspond to data taken while F2A filters were on. Pink/Cyan correspond to data taken with filters off.
  • Solid curves were taken with excitation ON and dashed curves were taken with excitation off.
  • We see good suppression of POS-> PIT coupling in MC2 and MC3. POS->YAw is minimally affected in all cases.
  • MC1 is clearly not doing good with the filters and probably needs readjustement. Something to do later in the future.
Attachment 1: IMC_F2A_Params_MC1.pdf
IMC_F2A_Params_MC1.pdf IMC_F2A_Params_MC1.pdf IMC_F2A_Params_MC1.pdf
Attachment 2: IMC_F2A_Params_MC2.pdf
IMC_F2A_Params_MC2.pdf IMC_F2A_Params_MC2.pdf IMC_F2A_Params_MC2.pdf
Attachment 3: IMC_F2A_Params_MC3.pdf
IMC_F2A_Params_MC3.pdf IMC_F2A_Params_MC3.pdf IMC_F2A_Params_MC3.pdf
Attachment 4: IMC_F2A_Foton.pdf
IMC_F2A_Foton.pdf IMC_F2A_Foton.pdf IMC_F2A_Foton.pdf
Attachment 5: MC1_POS2ANG_Filter_Test.pdf
MC1_POS2ANG_Filter_Test.pdf MC1_POS2ANG_Filter_Test.pdf MC1_POS2ANG_Filter_Test.pdf
Attachment 6: MC2_POS2ANG_Filter_Test.pdf
MC2_POS2ANG_Filter_Test.pdf MC2_POS2ANG_Filter_Test.pdf MC2_POS2ANG_Filter_Test.pdf
Attachment 7: MC3_POS2ANG_Filter_Test.pdf
MC3_POS2ANG_Filter_Test.pdf MC3_POS2ANG_Filter_Test.pdf MC3_POS2ANG_Filter_Test.pdf
  16088   Tue Apr 27 15:15:17 2021 Ian MacMillanUpdateCDSSUS simPlant model

The first version of the single filter plant is below. Jon and I went through compiling a model and running it on the docker (see this post)

We activated an early version of the plant model (from about 10 years ago) but this model was not designed to run on its own so we had to ground lots of unconnected pieces. the model compiled and ran so we moved on to the x1sus_single_plant model that I prepared.

This model is shown in the first attachment wasn't made to be run alone because it is technically a locked library (see the lock in the bottom left). It is supposed to be referenced by another file: x1sup.mdl (see the second attachment). This works great in the Simulink framework. I add the x1sus_single_plant model to the path and Matlab automatically attaches the two. but the docker does not seem to be able to combine the two. Starting the cymac it gives these errors:

cymac    | Can't find sus_single_plant.mdl; RCG_LIB_PATH=/opt/rtcds/userapps:/opt/rtcds/userapps/lib:/usr/share/advligorts/src/src/epics/simLink/:/usr/share/advligorts/src/src/epics/simLink/lib:/usr/share/advligorts/src/src/epics/simLink
cymac    | make[1]: *** [Makefile:30: x1sup] Error 2
cymac    | make: *** [Makefile:35: x1sup] Error 1

I have tried putting the x1sus_single_plant.mdl file everywhere as well as physically dragging the blocks that I need into the x1sup.mdl file but it always seems to throw an error. Basically, I want to combine them into one file that is not referencing anything other than the CDS library but I cant figure out how to combine them.

Okay but the next problem is the medm screen generation. When we had the original 2010 model running the sitemap did not include it. It included models that weren't even running before but not the model Jon and I had added. I think this is because the other models that were not running had medm screens made for them. I need to figure out how to generate those screens. I need to figure out how to use the tool Chris made to auto-generate medm screens from Simulink but I can't seem to figure it out. And honestly, it won't be much use to me until I can actually connect the plant block to its framework. One option is to just copy each piece over one by one. this will take forever but at this point, I am frustrated enough to try it. I'll try to give another update later tonight.

Attachment 1: x1sus_single_plant.pdf
x1sus_single_plant.pdf
Attachment 2: x1sup.pdf
x1sup.pdf
  16087   Tue Apr 27 10:05:28 2021 Anchal, PacoUpdateSUSMC1 and MC3 F2A Filters Tested

We extended the f2a filter implementation and diagnostics as summarized in 16086 to MC1 and MC3.


MC1

Attachment 1 shows the filters with Q=3, 7, 10. We diagnosed using Q=3.

Attachment 2 shows the test summary, exciting with broadband noise on the LSC_EXC and measuring the CSD to estimate the transfer functions.


MC3

Attachment 3 shows the filters with Q=3, 7, 10. We diagnosed using Q=3.

Attachment 4 shows the test summary, exciting with broadband noise on the LSC_EXC and measuring the CSD to estimate the transfer functions.


Our main observation (and difference) with respect to MC2 is the filters have relative success for the PIT cross-coupling and not so much for YAW. We already observed this when we tuned the DC output gains to compute the filters.

Attachment 1: IMC_F2A_Params_MC1.pdf
IMC_F2A_Params_MC1.pdf IMC_F2A_Params_MC1.pdf IMC_F2A_Params_MC1.pdf
Attachment 2: MC1_POStoAng_CrossCoupling.pdf
MC1_POStoAng_CrossCoupling.pdf MC1_POStoAng_CrossCoupling.pdf MC1_POStoAng_CrossCoupling.pdf MC1_POStoAng_CrossCoupling.pdf MC1_POStoAng_CrossCoupling.pdf MC1_POStoAng_CrossCoupling.pdf
Attachment 3: IMC_F2A_Params_MC3.pdf
IMC_F2A_Params_MC3.pdf IMC_F2A_Params_MC3.pdf IMC_F2A_Params_MC3.pdf
Attachment 4: MC3_POStoAng_CrossCoupling.pdf
MC3_POStoAng_CrossCoupling.pdf MC3_POStoAng_CrossCoupling.pdf MC3_POStoAng_CrossCoupling.pdf MC3_POStoAng_CrossCoupling.pdf MC3_POStoAng_CrossCoupling.pdf MC3_POStoAng_CrossCoupling.pdf
  16086   Mon Apr 26 18:55:39 2021 Anchal, PacoUpdateSUSMC2 F2A Filters Tested

Today we tested the F2A filters created from the DC gain values listed in 16066.


Filters:

  • For a DC gain G_{DC} required for balancing the coil at DC and f_0 being the resonance frequency of the mode (POS in this case), we calculate the filter using:
    \frac{1 + i \frac{f_z}{f Q} - \frac{f_z^2}{f^2}}{1 + \frac{f_0}{f} - \frac{f_0^2}{f^2}}where f_z = f_0 \sqrt{G_{DC}}.
  • Attachment 1 shows the motivation for choosing the resonant frequency in the formula above. It makes gain at DC as G_{DC} while keeping AC gain as 1.
  • Attachment 2 shows the transfer functions of the filters uploaded.
  • Filters are named Eg2CtQ3, Eg2CtQ7 and Eg2CtQ10 for Q=3,7,10 filters respectively. (Named for Eigenmode Basis to Cartesian Basis conversion filters, aka F2A filters).

Testing procedure:

  • We uploaded the new input matrix listed in 16066.
  • We then uploaded the coil output gains (AC gains) that are also listed in 16066.
  • Then we reduced the C1:IOO-WFS_GAIN to 0.05 (by a factor of 20).
    • Rana asked us to test the WFS sensors' impulse response to observe a minimum 10s decay to ensure that the UGF of WFS control loops is at or below 0.1 Hz.
    • We were unable to have any effect on this decay actually. We tried setting offsets without tramps in multiple places but whenever we were able to excite this loop, it will always damp down in about 5-6s regardless of the value of C1:IOO-WFS_GAIN.
    • So we moved on.
  • Then, with MC locked we took reference data with no excitation or filters uploaded. (dotted curves)
  • We took cross spectral density from C1:IOO-MC_F to C1:IOO-MC_TRANS_PIT_IN1, C1:IOO-MC_TRANS_YAW_IN1, C1:IOO-WFS1_PIT_IN1, C1:IOO-WFS1_PIT_IN1, C1:IOO-WFS2_PIT_IN1, and C1:IOO-WFS2_PIT_IN1.
  • We were also looking at the power spectral density of these channels.
  • Then using awggui (after the fix we did as in 16085), we added noise in C1:SUS-MC2_LSC_EXC as uniform noise between 0.05 Hz to 3.5 Hz with amplitude of 100 and gain of 100.
  • We took a set of data without switching on the filters to have a comparison later. (Dash-dort curves)
  • We then took data after switching on the filters. (Solid curves)

Next:

  • Tomorrow we'll repeat this for MC1 and MC3 if we get a favourable grade in our work here.
  • Even if not, we'll jsut conclude the suspension optimization work tomorrow morning and get into main interferometer.
Attachment 1: f2a.pdf
f2a.pdf
Attachment 2: IMC_F2A_Params_MC2.pdf
IMC_F2A_Params_MC2.pdf IMC_F2A_Params_MC2.pdf IMC_F2A_Params_MC2.pdf
Attachment 3: MC2_F2A_FilterChar_POS2Ang.pdf
MC2_F2A_FilterChar_POS2Ang.pdf MC2_F2A_FilterChar_POS2Ang.pdf MC2_F2A_FilterChar_POS2Ang.pdf MC2_F2A_FilterChar_POS2Ang.pdf MC2_F2A_FilterChar_POS2Ang.pdf MC2_F2A_FilterChar_POS2Ang.pdf
  16085   Mon Apr 26 18:52:52 2021 Anchal, PacoHowToComputer Scripts / Programsawg free slot

Today we had some trouble launching an excitation on C1:IOO-MC_LSC_EXC from awggui. The error read:

awgSetChannel: failed getIndexAWG C1:SUS-MC2_LSC_EXC ret=-3

What solved this was the following :

  1. launch the dtt command line interface
  2. Anchal remembers a slot number 37008
  3. We issue >> awg free 37008
  4. Slot freed, launch a new instance of awggui
ELOG V3.1.3-